For given analog signal xa(t) what is the maximum frequency of this signal and what is the Nyquist rate for this signal (choose the minimum one)

xa(t)=5 cos(310pi t) + 8 sin(420pi t)
Answer #1

211) - 500 (310174) + 8 sin (420114) < W = 3/01T W2 = 420 1T maximum frequne Wmax max (W7, W2) 42011 rad sec fmax = 42011 = 2